Pittsburgh Custom Products (PCP) purchased a new machine for ram-cambering large I beams. PCP expects to bend 55 beams at S3,000 per beam in each of the first 3 years, after which it expects to bend 100 beams per year at $2,700 per beam through year 11 If the companys minimum attractive rate of return is 16% per year, what is the present worth of the expected revenue? The present worth of the expected revenue is S
Answer #1

Please find below.......

we have to compute the present value by multiplying cash flow * discount factor
Computation of cash flow
i ii iii iv=ii*iii
Year revenue PVIF @ 16% present value
1 165000 =55*3000     0.8621      142,241.38
2 165000 =55*3000     0.7432      122,621.88
3 165000 =55*3000     0.6407      105,708.52
4 270000 =100*2700     0.5523      149,118.60
5 270000 =100*2700     0.4761      128,550.51
6 270000 =100*2700     0.4104      110,819.41
7 270000 =100*2700     0.3538        95,533.97
8 270000 =100*2700     0.3050        82,356.87
9 270000 =100*2700     0.2630        70,997.30
10 270000 =100*2700     0.2267        61,204.57
11 270000 =100*2700     0.1954        52,762.56
present value 1,121,915.58
Therefore present value = 1,121,915.58